A body starts from rest and with a uniform acceleration of 10 m/s2 for 5 seconds. During the next 10 seconds it moves with uniform velocity. The total distance travelled by the body is(1) 100 m(2) 125 m(3) 500 m(4) 625 m

Grade:11

1 Answers

Rohit Paliwal
6 years ago
Answer is option D 625 m When body starts for first 5 second with acceleration it covers 125mS=1/2at^2=1/2×10×5×5=125And for next 10 second the velocity is 50 therefore dosyance is 500mAnd 125m+ 500m =625m